# Zero-Shot Generalization during Instruction Tuning: Insights from Similarity and Granularity

### v1

@@ 3.2.1 Enhancing Zero-Shot Generalization through High-Similarity Data
Selection of Similarity Measures. To validate our hypothesis, we first have to define how to measure the similarity. To effectively permute all the training data, we aim to calculate the similarity distance between each training data point and the whole test set. Specifically, we investigate two main categories of similarity measures:
  • N-gram similarity: We respectively measure the similarity distance by calculating the KL divergence between the bigram word distributions of each training set data point and all the test set data.
  • Embedding similarity: We utilize all-MiniLM-L6-v2  [^3] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 from Sentence Transformer [36] to compute embeddings for each training and test data and then calculate the Cosine and Euclidean similarity distances, as detailed in Section B.3. Next, we refer to four classical distance calculation methods  [^4] https://en.wikipedia.org/wiki/Hierarchical_clustering, namely “max” (maximal distance), “min” (minimal distance), “avg” (unweighted average distance), and “centroid” (centroid distance), to represent the distance of each training set data point to all the test set data.
@@
Introduction of Weighted Similarity Distance. To raise the stability of computation and leverage the strengths from both Cosine-Avg and Cosine-Min, we consider performing a simple linear combination of both metrics for real similarity calculation. We denote our similarity distance measurement as Weighted Similarity Distance (WSD). Specifically, for each training data point $x_{i}$, we assume that the Average Distance and the Minimum Distance from $x_{i}$ to $\mathcal{D}_{\text{Test}}$ are $\text{AD}_{i}$ and $\text{MD}_{i}$ respectively. We define our WSD measure as follows:
$\text{WSD}_{i}=w_{\text{avg}}\text{AD}_{i}+w_{\text{min}}\text{MD}_{i}$ (1)
Empirically, we set $w_{\text{avg}}=1$, and $w_{\text{min}}$ to a scaling factor to ensure Cosine-Avg and Cosine-Min have the same magnitude, detailed in Equation 14. We investigate the impact of the WSD on zero-shot generalization during instruction tuning. Specifically, we examine three training permutations: Nearest First Training (NFT), Farthest First Training (FFT), and Random Training (RT) on the multi-task dataset Flan-mini. This setup allows us to differentiate between the nearest and farthest data points in terms of the temporal dimension of instruction tuning. By evaluating a series of fine-grained saved instruction-tuned checkpoints, we can compare the effects of the three settings and gain a clear understanding of the impact of WSD-based premutation on zero-shot generalization. More details are illustrated in Section B.1.
Settings. We utilize the Flan-mini dataset and randomly sample up to 20 instances for each training task. Each test task consists of at most five test data points to form a test set. We permute the training data based on the WSD measure to create the NFT, FFT, and RT settings. For each setting, we perform instruction tuning on the three training data permutations, resulting in a series of fine-grained checkpoints. We then calculate the average loss for each checkpoint on the test set containing various test tasks.
@@
Results. The earlier the model encounters data with high similarity to the test set, the more beneficial it is for zero-shot generalization. As shown in the left plot of Figure 3, we can observe that the NFT setting exhibits a rapid and low loss reduction, indicating better zero-shot generalization. In contrast, the FFT setting shows relatively poorer zero-shot generalization compared to the baseline RT setting.
3.2.2 Enhancing Zero-Shot Generalization through Fine-Grained Data
Traditional methods to enhance zero-shot generalization are mostly confined to the task level, focusing on task-pair transfer. However, the so-called “tasks” or “categories” are artificially defined and, from the perspective of LLMs, they are merely a collection of tokens or embedding representations. Therefore, different “tasks” or “categories” may still appear relatively similar to LLMs, while instances from the same task may exhibit profound differences. Thus, we propose and validate that:
{bclogo}
[…]
Settings. We use the Flan-mini dataset and randomly sample up to 20 instances for each training task. We employ two approaches to permute the training data: i) coarse-grained setting, where all instances under each training task are clustered and then shuffled. We define the embedding of a task as the average embedding of all instances under that task. ii) fine-grained setting, where all data instances are directly shuffled instead of first being clustered.
Results. Compared to the coarse-grained setting, the fine-grained setting is more beneficial for enhancing zero-shot generalization. As shown in the right plot of Figure 3, the loss curve for the fine-grained setting decreases more quickly and effectively, indicating that removing task framework constraints can further improve zero-shot generalization.
4 Understanding Zero-Shot Generalization
In previous experiments, we empirically apply Cosine-Avg and Cosine-Min, as well as the WSD measure obtained through weighting. However, these measures inherently possess shortcomings: i) Cosine-Avg cannot distinguish the variance within the test set. When the test set collapses to the point where all data points are the same, the distribution becomes a spike; when the test set data is sufficiently spread out, the distribution becomes uniform. In both cases, the Cosine-Avg distance from all test data points to a certain training data point remains consistent, indicating that it could not distinguish the role of each test data point. ii) Cosine-Min cannot holistically represent all test set data points. If a certain test data point is closer to all training data points than all other test data points, Cosine-Min only considers the distance to that test data point and does not take into account others, even if they are relatively far away.
Therefore, we seek an approach that can better distinguish and permute the training data for enhanced zero-shot generalization. To this end, we present the Test-centric Multi-turn Arrangement (TMA).
Algorithm 1 Test-centric Multi-turn Arrangement Method
1:Dataset $\mathcal{D}$ split into training set $\mathcal{D}_{\text{train}}$ and test set $\mathcal{D}_{\text{test}}$
2:Sub-training sets $\mathcal{D}_{\text{train}}^{i}$ for each round $i$
@@ 4 Understanding Zero-Shot Generalization
Formalization. We formalize TMA algorithm in Algorithm 1. This arrangement of the training data ensures that the embedding of each test data point is equally considered, thus taking into account all their characteristics. It avoids the limitations of Cosine-Avg, which disregards the distribution within the test data, or Cosine-Min, which only considers individual test data embeddings. A more detailed investigation of our arrangement method is provided in Section C.4.
4.1 Test-centric Multi-turn Arrangement Enhances Zero-Shot Generalization
@@
Data and Settings. We employ two types of datasets: i) datasets with task splits, such as Flan-mini [14], and ii) datasets without task splits, such as ShareGPT [46] and NoRobots [35]. Flan-mini consists of task-specific splits, while ShareGPT and NoRobots are general dialogue datasets. We arrange the training data by applying Algorithm 1 and examine the same three training permutations, namely NFT, FFT, and RT, which are consistent with the experimental setup described in Section 3.2.1. Specifically, NFT under this setting refers to the sequential training data order as returned by Algorithm 1, and FFT refers to its reverse. For detailed configurations, please refer to Section C.2.
Results. Using the Test-centric Multi-turn Arrangement to order training data from nearest to farthest significantly enhances zero-shot generalization. As illustrated in Figure 4, whether with the task split in Flan-mini (left) or without the task split in ShareGPT (middle) and NoRobots (right), the loss curve under the NFT setting decreases more rapidly while reaching a lower point, whereas the FFT setting results in the poorest performance. This further validates the effectiveness of our arrangement method. We also conduct an ablation study on the final distribution of $\mathcal{Q}_{\text{train}}$, as detailed in Section C.3. We reveal that accessing high-similarity data during instruction tuning can facilitate continual learning and further loss reduction.
4.2 Early-Selected Sub-Training Set Shows Higher Effectiveness
Figure 5: Averaged test loss on different turns of training data.
@@ 4.2 Early-Selected Sub-Training Set Shows Higher Effectiveness
Data and Settings. We apply Flan-mini as in the previous experiment. Specifically, we cluster the training data selected from rounds $[i,i+5)(i=0,5,10,15,20,25)$ based on TMA, and respectively perform instruction tuning. Then, for each instruction-tuned checkpoint, we calculate the average loss on the test set.
Results. The training data selected from the earliest rounds generally exhibits higher quality. As shown in Figure 5, we observe that from rounds $[0,5)$ to $[25,30)$, the decrease in loss becomes slower, and the minimum loss value tends to be higher. This indicates that our arrangement method could successfully distinguish high-quality training data and apply them early in tuning to enhance zero-shot generalization.
5 Related Work
In the course of deep learning’s development, several intriguing training dynamics have gradually been discovered, such as grokking [29], double descent [1], emergent abilities [49], and zero-shot generalization on unseen tasks. The first three phenomena have already received substantial explanations. For grokking, it happens from algorithm tasks [29] to a broader spectrum of realistic tasks [23], explained through the Slingshot Effects [42] or the competition between memorization and generalization circuits [44]. For double descent, Nakkiran et al. [27] proposed that double descent occurs both model-wise and epoch-wise, while Davies et al. [10] unified grokking and double descent through the learning speed and generalization ability of patterns. Regarding emergent abilities, Schaeffer et al.[38] suggested that the occurrence of emergent abilities is due to researchers using non-smooth metrics. Hu et al.[18] extended this idea by proposing an infinite resolution evaluation to predict the emergence of abilities, and Du et al. [12] attempted to understand emergent abilities from the perspective of loss. In the rapid development of LLMs, security issues have increasingly attracted the attention of researchers [8, 57]. Solving these security challenges relies on an in-depth understanding of training dynamics, which is crucial for comprehending the mechanisms of neural networks and contributes to building robust and interpretable LLMs. Research on these dynamics also offers valuable insights for investigating zero-shot generalization, especially the work on emergent abilities. We suggest that model-wise emergent abilities may be analogous to step-wise zero-shot generalization.
LLMs have been proven capable of zero-shot generalization across a variety of downstream tasks [2], and instruction tuning has emerged as the most effective method to achieve this [5, 37, 48]. Instruction-tuned LLMs possess strong capabilities in instruction following and logical reasoning. Moreover, enhancing their understanding of the implicit intentions behind instructions [31] makes it possible to develop these models into powerful agents [30, 34, 32, 33]. The zero-shot generalization phenomenon resulting from instruction tuning is crucial for building general LLMs. This means that models trained on certain tasks can generalize well to unseen tasks. Multi-task datasets designed for this purpose have continuously iterated in quality, quantity, and diversity, and numerous studies have explored how zero-shot generalization occurs during the instruction tuning process. Sanh et al. [37] constructed P3 using explicit multitask learning, demonstrating that explicit task prompt templates can promote zero-shot generalization. Wang et al. [47] created the Super Natural Instructions V2 (NIV2) dataset, which comprises over 1600 task types, and empirically showed that more observed tasks, an adequate number of training instances, and larger models improve generalization. Meta introduced OPT-IML [19], investigating the impacts of dataset scale and diversity, different task sampling strategies, and the presence of demonstrations on generalization. Subsequently, Chung et al. [24] proposed the Flan Collection, which encompasses up to 1836 tasks, and pointed out that scaling the number of tasks and model size, as well as incorporating chain-of-thought data, can dramatically improve performance on unseen tasks.
In addition, a line of research focuses on understanding the relationships in task-pair transfer [40, 45, 52], suggesting that not all tasks contribute positively to zero-shot generalization; some tasks may even result in negative transfer effects [21, 26, 56]. However, a significant limitation of the aforementioned works is that, whether training on one task and then evaluating on another [21, 56], or simply calculating intermediate task [45] or instruction [22] transfer scores, all these efforts to select the most informative tasks to promote zero-shot generalization are confined within the “task” framework. This approach is based on a premise: the human-defined “tasks” and even “categories” are sufficiently reasonable. This is precisely the issue our study strives to address: breaking free from the task-level framework to explore “generalization” at a more fundamental level.
@@ 6 Conclusion
Our research sheds light on the mechanism underlying zero-shot generalization during instruction tuning, moving beyond the conventional task-level analysis to a more data-centric perspective. By demonstrating that zero-shot generalization occurs early during instruction tuning and is significantly influenced by data similarity and granularity, we provide a new understanding of how instruction tuning brings up zero-shot generalization. Our proposed Test-centric Multi-turn Arrangement further illustrates the importance of accessing high-similarity data early in the training process to facilitate continual learning and loss reduction. For future work, we suggest exploring the quantitative relationship between similarity distance and loss. Specifically, investigating whether similarity distance can predict a model’s generalization performance on new data could further help the optimization of instruction tuning. We hope our findings will pave the way for developing more aligned and robust LLMs, enhancing their ability to generalize effectively in diverse applications.
@@ A.1 Data and Setting
We utilized three datasets: Super Natural Instructions V2 [47], Public Pool of Prompts [37] and Flan-mini [14]. Here, we provide a detailed overview of each dataset.
@@ A.3 Hyper-Parameter Details
For instruction tuning, we present some key hyper-parameters related to instruction tuning in Table 2. Additionally, we utilize the model-center framework [25] to conduct full-parameter instruction tuning of LLaMA-2-7B on two 80GB A800s and dynamically adjust the loss scale based on the changing training loss to prevent underflow. All of our instruction tuning experiments utilize these hyper-parameters consistently.
@@ A.4 Evaluation Details
  • ROUGE-1: ROUGE-1 measures the comprehensiveness of the generated summary by calculating the overlap between words in the generated summary and words in the reference summary:
    ROUGE-1=Number of overlapping unigramsTotal number of unigrams in reference $\text{ROUGE}\textnormal{-}1=\frac{\text{Number of overlapping unigrams}}{\text{Total number of unigrams in reference summary}}$ (2)
  • ROUGE-L: ROUGE-L is based on the idea of Longest Common Subsequence (LCS). By measuring the length $m$ of the reference summary $X$ and the length $n$ of the generated summary $Y$, the ROUGE-L score is calculated as follows:
@@
    $\text{EM}=\begin{cases}1&\text{if the model's prediction exactly matches the true answer}\\ 0&\text{otherwise}\end{cases}$ (6)
@@ A.5 Discussion for Metrics
In previous experiments, we have discovered that zero-shot generalization might occur early in the instruction tuning process based on the ROUGE series, Exact-Match, and RM score. However, these metrics may not be suitable for measuring generalization effectively. First, for the ROUGE series, ROUGE-1 refers to the overlap of unigrams, and ROUGE-L is based on the longest common subsequence. Both metrics are limited to surface-level matching, primarily relying on lexical overlap between model outputs and labels, to the extent that capturing semantic similarity or a deeper understanding of the content conveyed in the sentences becomes challenging [13, 15]. Outputs and labels with different wordings but similar meanings may receive low ROUGE scores.
While the reliability of ROUGE series scores is questionable, metrics like Exact-Match are nonlinear, and previous research [38] has shown that nonlinear metrics are prone to observing emergent abilities. Although emergence is a model-wise phenomenon, if we adopt such nonlinear metrics step-wise, i.e., along the training timeline, we might also observe step-wise “emergence” so-called generalization phenomena. This might lead to misjudgments regarding the timing of zero-shot generalization. Therefore, we need to address this issue.
@@
Acknowledging that the Reward Model (RM) is trained on preference data, it is inevitable that there will be a certain loss of ability when generalizing to out-of-distribution (OOD) datasets [39]. Consequently, scoring on different datasets may not be precise. As shown in Figure 1, RM scores for NIV2 are notably higher than those for the other two datasets, indicating a bias. Further, we compare the reward distribution with respect to the correctness of the model response respectively on the three datasets. The large overlap under both curves in Figure 6 indicates that RM could not well distinguish between responses of lower or higher quality. This further highlights the inappropriateness of using RM score to reflect zero-shot generalization.

### v2

@@ 4.2.1 Effect of High-Similarity Data
Selection of Similarity Measures. To validate our hypothesis, we first define what similarity measures to apply. Based on our setting and previous works, we investigate two main categories of similarity measures:
  • N-gram similarity: We respectively measure the similarity distance by calculating the KL divergence between the bigram word distributions of the training set data and test set data.
  • Embedding similarity: We utilize all-MiniLM-L6-v2  [^3] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 from Sentence Transformer (Reimers and Gurevych, 2019) to compute embeddings for each training and test data and then calculate the Cosine and Euclidean similarity distances, as detailed in Section B.3. Next, we refer to four classical distance calculation methods  [^4] https://en.wikipedia.org/wiki/Hierarchical_clustering, namely “max” (maximal distance), “min” (minimal distance), “avg” (unweighted average distance), and “centroid” (centroid distance), to represent the distance from the training set data to test set data.
@@
Data and Settings. We utilize the Flan-mini dataset and randomly sample up to 20 instances for each training task. Each test task consists of at most five test data points to form a test set. For each training data point $x_{i}$, we calculate the similarity measure score between $x_{i}$ and the whole test set $\mathcal{D}_{\text{Test}}$. We arrange the training data based on this score. Specifically, we examine three training arrangements: Nearest First Training (NFT), Farthest First Training (FFT), and Random Training (RT), as shown in LABEL:fig:exp_settings_2. This setup allows us to differentiate between the nearest and farthest data points in terms of the temporal dimension of instruction tuning. We perform instruction tuning on the three training data arrangements, resulting in a series of fine-grained checkpoints. And then calculate the average loss for each checkpoint on the test set containing various test tasks.
Results. The earlier the model encounters data with high similarity to the test set, the more beneficial it is for zero-shot generalization. As shown in the left plot of Figure 3, we can observe that the NFT setting exhibits a rapid and low loss reduction, while the FFT setting shows relatively poorer zero-shot generalization compared to the baseline RT setting.
4.2.2 Effect of Fine-Grained Data
[…]
Data and Settings. We use the Flan-mini dataset and randomly sample up to 20 instances for each training task. We employ two approaches to arrange the training data: i) coarse-grained setting, where all instances under each training task are clustered. We define the embedding of a task as the average embedding of all instances under that task and arrange the clusters based on NFT, as shown in LABEL:fig:intro. ii) fine-grained setting, where all data instances are directly arranged basen on NFT instead of being clustered first.
Results. Compared to the coarse-grained setting, the fine-grained setting is more beneficial for improving zero-shot generalization. As shown in the right plot of Figure 3, the loss curve for the fine-grained setting decreases more quickly and effectively, indicating that removing task framework constraints can further improve generalization.
@@ 4.2.2 Effect of Fine-Grained Data
5 Test-centric Multi-turn Arrangement
In earlier experiments, we show that training data arrangement based on combined similarity measures affects zero-shot generalization. However, conventional similarity measures have limitations: i) Cosine-Avg fails to capture variance within the test set, as it remains unchanged whether the test set is highly clustered or uniformly spread. ii) Cosine-Min offers a limited perspective, focusing solely on the nearest test point while ignoring the overall test distribution.
Therefore, we seek an approach that can better distinguish and arrange the training data for analyzing zero-shot generalization. To this end, we present the Test-centric Multi-turn Arrangement (TMA) framework.
Algorithm 1 Test-centric Multi-turn Arrangement
1:Training set $\mathcal{D}_{\text{train}}$ and test set $\mathcal{D}_{\text{test}}$
2:Sub-training sets $\mathcal{D}_{\text{train}}^{i}$ for each turn $i$
@@ 5 Test-centric Multi-turn Arrangement
Formalization. We formalize the TMA framework in Algorithm 1. In this manner, we progressively construct the training set by selecting subsets that are decreasingly similar to the test data. These sub-training sets can then be arranged in either a Nearest-First Training (NFT) or Farthest-First Training (FFT) manner for further experiments. This arrangement of the training data ensures that the embedding of each test data point is equally considered, thus taking into account all their characteristics. A more detailed investigation of our arrangement method is provided in Section C.3.
5.1 TMA Improves Zero-Shot Generalization
@@
Data and Settings. We employ two types of datasets: i) datasets with task splits, such as Flan-mini (Ghosal et al., 2023), and ii) datasets without task splits, such as ShareGPT (Wang et al., 2023) and NoRobots (Rajani et al., 2023). Flan-mini consists of task-specific splits, while ShareGPT and NoRobots are general dialogue datasets. We arrange the training data by applying Algorithm 1 and examine the same three training arrangements, namely NFT, FFT, and RT, which are consistent with the experimental setup in Section 4.2.1. Specifically, NFT under this setting refers to the sequential training data order as returned by Algorithm 1, and FFT refers to its reverse. For detailed configurations, please refer to Section C.2.
Results. Using our proposed TMA to arrange training data from nearest to farthest improves zero-shot generalization. As illustrated in Figure 4, whether with the task split in Flan-mini (left) or without the task split in ShareGPT (middle) and NoRobots (right), the loss curve under the NFT setting decreases more rapidly while reaching a lower point, whereas the FFT setting results in the poorest performance. This validates the effectiveness of our arrangement method TMA.
5.2 Ablation Study
@@
Data and Settings. For Flan-mini, we randomly select 225 tasks as the test set and use all data from the remaining tasks as the training set. We apply TMA to the entire training set, comprising over 1 million instances. Training data from different turns, denoted as $\mathcal{D}_{\text{train}}^{i}$ ($i\in[1,N]$, where $N$ is the total number of turns), are organized under five strategies, with $M$ representing the desired number of samples:
  • NFT (Nearest First Training): Data is sequentially selected from $i=1$ to $i=N$ until reaching $M/2$ samples ($\mathcal{D}_{\text{train1}}$), then from $i=N$ to $i=1$ for another $M/2$ samples ($\mathcal{D}_{\text{train2}}$). The two subsets are merged, maintaining nearest-to-farthest ordering.
  • FFT (Farthest First Training): Similar to NFT, but the merged data is ordered from farthest to nearest.
  • MAX: Data is sequentially selected from $i=N$ to $i=1$ until $M$ samples are accumulated.
  • MIN: Data is sequentially selected from $i=1$ to $i=N$ until $M$ samples are accumulated.
Results. Early exposure to similar training data aids generalization while accessing high-similarity data during instruction tuning facilitates continual learning and further loss reduction. From LABEL:fig:macro_research, we observe:
  • NFT and MIN: Loss curves exhibit similar patterns, indicating that early exposure to training data resembling the test set benefits generalization the most.
  • FFT and MAX: Loss curves diverge midway (around 950 steps), as FFT encounters high-similarity training data, reducing loss further. This highlights the advantage of high-similarity data during instruction tuning.
  • RT: Positioned between NFT and FFT, RT serves as a baseline, demonstrating intermediate performance.
@@ 6 Conclusion
Our research sheds light on the mechanism underlying zero-shot generalization during instruction tuning, moving beyond the conventional task-level analysis to a more data-centric and fine-grained perspective. By demonstrating that zero-shot generalization occurs early during instruction tuning and is significantly influenced by data similarity and granularity, we provide a new understanding of how instruction tuning brings up zero-shot generalization. The Test-centric Multi-turn Arrangement framework further illustrates the importance of accessing high-similarity data early in the training process to facilitate continual learning and loss reduction. For future work, we suggest exploring the quantitative relationship between similarity distance and loss. Specifically, investigating whether similarity distance can predict a model’s generalization performance on new data could further help the optimization of instruction tuning. We hope our findings will pave the way for developing more aligned and robust LLMs.
Limitations
Although our research has made significant progress by discovering that zero-shot generalization occurs in the early stage of instruction tuning and proposing various similarity distance measures to explore their impact on zero-shot generalization, we acknowledge that our study is far from perfect. Firstly, conducting a single experiment can be costly due to storage space requirements and computational resource limitations, so we only conducted limited explorations on LLaMA-2-7B with a few runs, which may introduce biases in our conclusions. Secondly, the similarity distance measures we proposed may not have a strong theoretical foundation and can only serve as supplements to existing measures. Lastly, we chose loss as the metric for zero-shot generalization instead of traditional task-level evaluations often seen in benchmarks with objective metrics. This is because we believe that traditional task-level generalization has certain limitations, as different tasks or categories may still appear relatively similar to LLMs, while instances from the same task may exhibit profound differences. However, this viewpoint still requires further validation. We hope future works can address these limitations.
Broader Impacts
@@ A.1 Data and Setting
We utilized three datasets: Super Natural Instructions V2 Wang et al. (2022), Public Pool of Prompts Sanh et al. (2021) and Flan-mini Ghosal et al. (2023). Here, we provide a detailed overview of each dataset.
@@ A.3 Hyper-Parameter Details
For instruction tuning, we present some key hyper-parameters related to instruction tuning in Table 2. Additionally, we utilize the model-center framework (modelcenter, 2023) to conduct full-parameter instruction tuning of LLaMA-2-7B on two 80GB A800s for 8 hours and dynamically adjust the loss scale based on the changing training loss to prevent underflow. All of our instruction tuning experiments utilize these hyper-parameters consistently.
@@ A.4 Evaluation Details
  • ROUGE-1: ROUGE-1 measures the comprehensiveness of the generated summary by calculating the overlap between words in the generated summary and words in the reference summary. Let $n_{o}$ be the number of overlapping unigrams and $n_{r}$ be the total number of unigrams in the reference summary. Then:
    $\text{ROUGE}\textnormal{-}1=\frac{n_{o}}{n_{r}}$ (1)
  • ROUGE-L: ROUGE-L is based on the idea of Longest Common Subsequence (LCS) and we use rouge-score library  [^7] https://github.com/google-research/google-research/tree/master/rouge for implementation. By measuring the length $m$ of the reference summary $X$ and the length $n$ of the generated summary $Y$, the ROUGE-L score is calculated as follows:
@@
    $\text{EM}=\begin{cases}1&\text{if prediction = reference}\\ 0&\text{otherwise}\end{cases}$ (5)
@@ A.5 Discussion for Metrics
In previous experiments, we have discovered that zero-shot generalization might occur early in the instruction tuning process based on the ROUGE series, Exact-Match, and RM score. However, these metrics may not be suitable for measuring generalization effectively. First, for the ROUGE series, ROUGE-1 refers to the overlap of unigrams, and ROUGE-L is based on the longest common subsequence. Both metrics are limited to surface-level matching, primarily relying on lexical overlap between model outputs and labels, to the extent that capturing semantic similarity or a deeper understanding of the content conveyed in the sentences becomes challenging Ganesan (2018); Grusky (2023). Outputs and labels with different wordings but similar meanings may receive low ROUGE scores.
While the reliability of ROUGE series scores is questionable, metrics like Exact-Match are nonlinear, and previous research Schaeffer et al. (2024) has shown that nonlinear metrics are prone to observing emergent abilities. Although emergence is a model-wise phenomenon, if we adopt such nonlinear metrics step-wise, i.e., along the training timeline, we might also observe step-wise “emergence” so-called generalization phenomena. This might lead to misjudgments regarding the timing of zero-shot generalization. Therefore, we need to address this issue.
@@
Acknowledging that the Reward Model (RM) is trained on preference data, it is inevitable that there will be a certain loss of ability when generalizing to out-of-distribution (OOD) datasets Singhal et al. (2024). Consequently, scoring on different datasets may not be precise. As shown in Figure 1, RM scores for NIV2 are notably higher than those for the other two datasets, indicating a bias. Further, we compare the reward distribution with respect to the correctness of the model response respectively on the three datasets. The large overlap under both curves in Figure 5 indicates that RM could not well distinguish between responses of lower or higher quality. This further highlights the inappropriateness of using RM score to reflect zero-shot generalization.